# TextCenGen: Attention-Guided Text-Centric Background Adaptation for Text-to-Image Generation

### v3

@@ 4.2 Comparison with Existing Methods
To evaluate model performance, we used metrics assessing various aspects of the generated images. The CLIP Score Hessel et al. (2021); Huang et al. (2021); Radford et al. (2021); cli (2022) measured semantic fidelity, ensuring images align with textual descriptions. The total variation (TV) loss Rudin and Osher (1994) assessed the visual coherence and smoothness of the background in relation to text region $R$, which is crucial for harmonious compositions. Saliency Map Intersection Over Union (IOU) quantified focus and clarity around text areas. We propose the Visual-Textual Concordance Metric (VTCM) that combines a global metric increasing with value (CLIP Score) and local metrics that benefit from lower values (Saliency IOU and TV Loss) within $R$. The VTCM formula is:
@@
Our quantitative analysis in Table 1 shows that our method, TextCenGen, outperforms native stable diffusion 1.5, especially in background smoothness and saliency harmony, even surpassing the latest Dalle-3. Notably, despite DiffText achieving the best results in terms of background significance as represented by the Saliency IOU and the region smoothness (TV loss), its CLIP score, which measures semantic similarity, significantly lags behind other methods. Scene text rendering often blurs text locations, unlike our approach of moving main objects to create space. Our method results in more natural and harmonious text layouts with highest VTCM.
@@
Our qualitative analysis, shown in  Figure 6, involved a comparison across different models using prompts and positions within the same quadrant. Dall-E 3, which relies solely on text inputs, exhibited significant variability and could not consistently clear the necessary areas for text. DiffText was effective in creating large blank spaces but often failed to fulfill the prompt accurately. For example, in the cases of ’bicycle’ and ’canny tree’, the central concepts were missing from the images. This probably occurs because the method reduces attention outside the region $R$, leading to a mismatch between the result image and the intended semantics of the prompt. Introducing text to images was tricky, as TV Loss showed, but TextCenGen kept object detail and background quality well. It shows that our force-directed method effectively balances text and visuals in images.
4.3 Ablation Study
Figure 7 presents the results of our ablation study. This study was designed to elucidate the individual contributions of the various components embedded within our proposed model. As part of our comprehensive analysis, we focused on assessing the two key factors: (1) the impact of the Force-Directed Cross-Attention Guidance, (2) the effectiveness of the Spatial Excluding Cross-Attention Constraint.
Impact of Force-Directed Cross-Attention Guidance. Our model’s force-directed feature is key to gently shifting where objects are placed. Without this, we might just bluntly edit the cross-attention map, which could mess up important parts of the picture. This part of our model helps us make sure we don’t ruin the image structure by harshly removing attention map from areas.
Effects of Spatial Excluding Cross-Attention Constraint. Despite successful relocation of conflict object-related tokens, spaces left behind may not inherently lead to a well-blended transition. Our experimental results underscore that the integration of the Spatial Excluding Cross-Attention Constraint greatly improve the smoothness of the remaining image sections.
5 Conclusion
@@
Our approach has certain limitations. The force-directed cross-attention guidance, assuming convexity and centering on object centroids, may not be suitable for non-convex shapes. This could lead to reduced size or fragmentation of objects. Future work will address these challenges to improve the quality of the output images. Additionally, we observed the generation of unintended objects in blank areas not mentioned in the original prompts. Future work could use BLIP-2 Li et al. (2023b) to update image prompts and detect new conflict objects.

### v4

@@ Introduction
To evaluate model performance, we used metrics assessing various aspects of the generated images. We proposed the CLIP Score Loss to evaluate the reduction in prompt semantic alignment for the training-free method compared to the original diffusion model. The CLIP Score (Hessel et al. 2021; Huang et al. 2021; Radford et al. 2021) measured semantic fidelity, ensuring images align with textual descriptions. The total variation (TV) loss (Rudin and Osher 1994; Jiang, Li, and Wang 2021) assessed the visual coherence and smoothness of the background in relation to text region $R$, crucial for harmonious compositions. Saliency Map Intersection Over Union (IOU) quantified focus and clarity around text areas. We propose the Visual-Textual Concordance Metric (VTCM) that combines a global metric increasing with value (CLIP Score) and local metrics that benefit from lower values (Saliency IOU and TV Loss) within $R$. The VTCM formula is:
@@
Our quantitative analysis in Table 1 and Table 2 shows that TextCenGen minimizes semantic loss while maintaining smoothness and saliency harmony within region R, even surpassing the latest Dalle-3. Scene text rendering and attention reduction in Desigen often focus on local attention, neglecting global semantics. Especially in Text in Object situation (see Figure 3), this weakens the ability to create whitespace. Our approach moves main objects away, then reduces attention to create space, resulting in more natural and harmonious text layouts with the highest VTCM.
MLLM-as-Judge ELO Ranking.
Following rising trend of multimodal large language model (MLLM) as judge methods (Chen et al. 2024; Wu et al. 2024), we present the MLLM-as-Judge ELO ranking for design appeal across different datasets in Table 3. The results demonstrate that the structured output from the GPT-4o with structured output provides consistent ratings, indicating that our method, along with Anytext and Dall-E, shows advantages in terms of design appeal. Interestingly, Dall-E excels in the LLM dataset that contains a higher proportion of natural landscapes. For detailed information on the prompts and evaluation methodology, please refer to the appendix.
Rank Method DDB P2P LLM
1 TextCenGen 702.21 752.85 122.96
2 Anytext 279.56 329.32 -89.68
3 Dall-E -17.32 -39.05 78.87
4 Desigen -322.63 -291.80 -7.24
5 SD1.5 -629.33 -738.83 -92.40
Table 3: Method ELO design appealing rankings across three datasets.
@@
Our qualitative analysis, shown in Figure 4, involved a comparison across different models using prompts and positions within the same quadrant. Dall-E 3, which relies solely on text inputs, exhibited significant variability and could not consistently clear the necessary areas for text. Desigen reduces attention in region $R$, but this method is not always effective without pretrained graphic design-specific model weights, especially when region $R$ is within a main object, such as in the bicycle example. Introducing text to images was tricky, as TV Loss showed, but TextCenGen kept object detail and background quality well. It shows that our force-directed method effectively balances text and visuals in images.
User Study
To understand the importance of the text-friendliness issue and explore users’ subjective perceptions of the T2I Model’s results, we conducted a user study with 111 participants. We used the qualitative results (see Figure 4) to develop a questionnaire. Figure 5 illustrates that the prompt-image alignment and aesthetics of our results were well-received in human perception.
Ablation Study
Table 4 presents the results of our ablation study. As part of our comprehensive analysis, we evaluated the two main contributors: (1) the impact of the force-directed component and (2) the effectiveness of the Spatial Excluding Cross-Attention Constraint.
Impact of Force-Directed Cross-Attention Guidance.
Our model’s force-directed feature is key to gently shifting where objects are placed. Without this, we might just bluntly edit the cross-attention map, which could mess up important parts of the picture. This part of our model helps us make sure we don’t ruin the image structure by harshly removing attention map from areas.
Effects of Spatial Excluding Cross-Attention Constraint.
Despite successful relocation of conflict object-related tokens, spaces left behind may not inherently lead to a well-blended transition. Our experimental results underscore that the integration of the Spatial Excluding Cross-Attention Constraint improve the smoothness of the remaining image sections.
Figure 5: User study of task importance and result evaluation. The left shows user perceptions of task encounter frequency (a1) and importance (a2), rating 5 as the highest. The right side details user rankings (b2) and the average times our results were chosen in a multiple-selection scenario (b1). The y-axis in b2 represents rankings from 1 to 5, demonstrating significant mean differences ($\alpha$=0.05) across three standards for our method.
w/o Both w/o FDG w/o SEC Ours
CLIP Score Loss $\downarrow$ - 2.2 1.71 0.32
TV Loss $\downarrow$ 14.39 12.44 12.76 13.67
Saliency IOU $\downarrow$ 30.32 28.56 28.61 28.34
VTCM $\uparrow$ 2.93 3.03 3.05 3.07
Table 4: Ablation study result. We examine TextCenGen with or without the implementation of Force-Directed Cross-Attention Guidance (FDG) and Spatial Excluding Cross-Attention Constraint (SEC). The CLIP Score Loss of w/o both indicates the use of the original SD-1.5 without our training-free method, resulting in no loss (-).
Conclusion
@@
Our approach has certain limitations. The force-directed cross-attention guidance, assuming convexity and centering on object centroids, may not be suitable for non-convex shapes. It may lead to reduced size or fragmentation of objects. Future work will address these challenges to improve the quality of the output images. Additionally, we observed the generation of unintended objects in blank areas not mentioned in the original prompts.